OpenStudy (anonymous):

simplify if possible log0.01

jimthompson5910 (jim_thompson5910):

Hints 1) 0.01 = 1/100 = 1/(10^2) = 10^(-2), so in short, 0.01 = 10^(-2) 2) log(x^y) = y*log(x) 3) log(10) = 1 where the base of the log is 10

OpenStudy (anonymous):

I was thinking it was log 0.01 = 10^-1 because there was only one zero after the decimal

jimthompson5910 (jim_thompson5910):

0.1 = 10^(-1)

OpenStudy (anonymous):

oh ok that makes sense!

jimthompson5910 (jim_thompson5910):

log(0.01) = log(10^(-2)) log(0.01) = -2*log(10) ... use the rule in hint #2 above log(0.01) = -2*1 ... use the rule in hint #3 above log(0.01) = -2 and I'm sure you probably figured it out by now

OpenStudy (anonymous):

If I were to write it as log 0.01 = 10^-2 would that still be correct?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

0.01 = 10^(-2), that's true but log(0.01) = 10^(-2) is false

OpenStudy (anonymous):

ok. for this next one its asking me to use a calculator to approximate the logarithm to four decimals (round to four decimal places as needed) log 2.68

OpenStudy (anonymous):

I dont have a scientific calculator.

jimthompson5910 (jim_thompson5910):

well you can use google as a calculator typing in log(2.68) will give you this https://www.google.com/search?hl=&q=log%282.68%29+&sourceid=navclient-ff&rlz=1B3GGLL_enUS420US420&ie=UTF-8

jimthompson5910 (jim_thompson5910):

make sure you round to 4 places

OpenStudy (anonymous):

so it would be 0.4281

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

sure thing